OpenStudy (anonymous):

which of the following are NOT trigonometric identities, Check all that apply a.) sec^(2) x - tan^(2) x=1 b) sec^(2) x+csc^(2) x=1 c) sin^(2) x + cos^(2) x=1 d) tan^(2) x + sec^(2) x =1

OpenStudy (anonymous):

i was thinking of B

OpenStudy (anonymous):

just b

OpenStudy (anonymous):

a) and c) only are trignometric identities

OpenStudy (zehanz):

Check them one by one: a.\[\sec^{2} x - \tan^{2} x=\frac{ 1 }{ \cos^2x }-\frac{ \sin^2x }{\cos^2x }=\frac{ 1-\sin^2x }{ \cos^2x }=\frac{ \cos^2x }{ \cos^2x }=1\], so a. is true...

OpenStudy (anonymous):

d is trigonometric too dude.

OpenStudy (anonymous):

@darth How?